# A Survey of Deep Learning Library Testing Methods

### v2

@@ 1. Introduction
However, existing surveys on the DL library testing are limited. We have compared related surveys in Table 1 and observed that existing works mainly focus on testing and repairing DL models or software built upon the models (Braiek and Khomh, 2020; Gezici and Tarhan, 2022; Martínez-Fernández et al., 2022; Zhang et al., 2022; Hu et al., 2023) (italic columns in Table 1). Even though some works involve testing DL libraries, they typically only cover certain components of various libraries and fail to provide a detailed introduction, analysis, and discussion across testing methods for different DL library components. For example, Zhang et al. (Zhang et al., 2020a) and Ji et al. (Ji et al., 2023) discussed testing methods for DL frameworks, while Li et al. (Li et al., 2020b) focused on the workflow and performance of DL compilers. To fill this gap, this paper focuses on the bugs that affect the functionality correctness and efficiency of DL libraries and comprehensively summarizes corresponding testing methods for three DL library components, namely DL frameworks, DL compilers, and DL hardware libraries. This paper further analyzes the advantages and limitations of existing methods and delves into the challenges and future research opportunities in DL library testing. Our goal is to provide a set of practical findings to promote the development of DL library testing research, thereby ensuring the security and reliability of DL systems built on these libraries in real-world scenarios. Note that the scope of this paper includes the testing methods on three DL library components, but does not include the DL model and DL program testing. The main contribution of this work can be summarized as follows:
  • We present the first comprehensive and detailed survey of testing methods for DL libraries, filling the gap of existing DL testing survey work that focuses heavily on DL models and lacks insight into the various types of bugs on DL libraries.
@@ 2.2. DL Library Components
DL libraries implement the abstract DL model based on the program and perform specific operations and optimizations on the underlying hardware to obtain computation results of training of DL model training and inference. Executing the DL program and building a DL model on the DL underlying libraries mainly involves three components, namely DL framework, DL compiler, and DL hardware library. The three components are closely related, cooperate with each other, and together form the DL workflow, as shown in Fig. 1. The code in dotted boxes presents the demo inputs of each component. Developers first need to call the DL framework APIs to construct a DL program. Overall, the DL framework executes the DL program and constructs the corresponding DL model. Then the abstract model will be passed to the DL compiler, and the DL compiler translates the input model into intermediate representations (IRs), optimizes it according to the target hardware, and outputs optimized operators and code. Finally, the DL hardware library (e.g., cuDNN) accepts the output of the compiler and maps calculations to the DL hardware to perform calculations and obtain the results.
@@ 3. Research Method
RQ 4.2 What are the challenges in existing DL library testing
@@ 5.1.2. Differential Testing on DL Framework
Model-level Differential Testing. The model-level differential testing usually leverages the different results of a widely-used DL model (e.g., ResNet-50) on different platforms or frameworks to detect bugs (Wang et al., 2020b; Li et al., 2022a). CRADLE (Pham et al., 2019) is one of the first tools to detect and identify bugs based on the concept of differential testing. Based on Keras (Ketkar and Ketkar, 2017) which can build and train models on different DL frameworks as backends, CRADLE conducted differential testing on three frameworks(i.e., TensorFlow, CNTK, and Theano) and finally detected 12 bugs. Fig. 7 shows two trigger figures of the inconsistencies that cause one model to have different prediction results and accuracy on different DL frameworks. CRADLE compared the model layer outputs between multiple DL frameworks and detected such inconsistencies. However, the inconsistent outputs of one layer may further lead to inconsistencies in subsequent layers, therefore this localization method is prone to false positives (FPs) and false negatives (FNs). To break these limitations, Guo et al. (Guo et al., 2020) mutated the parameters of model layers to explore more DL model behaviors and leveraged a causal-testing-based technique to localize buggy layers and reduce the FPs. Although the above testing methods effectively detected bugs in popular DL frameworks like TensorFlow, Theano, and CNTK, they only focused on the model inference process and could not detect bugs in the backward propagation stage. To fill the research gap, Gu et al. (Gu et al., 2022a) proposed Muffin, which creatively generated structure information and layer information of the DL model to thoroughly explore possible abnormal behaviors of the model. Then it conducted differential testing between DL frameworks for model training and inference processes and finally detected 39 new bugs. Although model-level differential testing methods obtain outstanding test results, they still have limitations in practice. The major limitation is that they can only explore limited DL model-related APIs and implementations in the framework For example, existing research (Wei et al., 2022) indicated that LEMON only covered 35 TensorFlow APIs. As a result, these methods cannot comprehensively test the entire framework (including APIs that can be used to develop ML algorithms), limiting the effectiveness of model-level differential testing methods.
@@ 5.1.4. Metamorphic Testing on DL Framework
Metamorphic testing has been widely used to detect the problems in DL models and systems (Zhang et al., 2021c; Arrieta, 2022). Despite the success of metamorphic testing in DL testing, researchers have paid limited attention to the metamorphic testing of DL frameworks. Ding et al. (Ding et al., 2017) constructed 11 MRS and validated AlexNet on the Caffe DL Framework. Although they were the first to try to design MRs to validate a DL framework, they focused on the DL model accuracy in tests and paid less attention to DL framework bugs. To address the limitation of localizing DL framework bugs at the parameter and tensor level, Chen et al. (Chen et al., 2024) proposed a testing framework with 18 MRs to test 10 common DL operators. However, their manually designed MRs only cover popular DL operators like Conv2d and lack support for the thousands of other DL framework APIs.
@@ 5.1.5. Other Testing on DL Framework
Comparison Experiment and Results. To compare the effectiveness of testing methods based on different testing techniques and principles, we conducted experiment with five state-of-the-art and representative open-source testing methods, covering three popular testing techniques and both model-level and API-level methods, namely Muffin (Gu et al., 2022a), FreeFuzz (Wei et al., 2022), DocTer with conforming (CI) and violating inputs (VI) (Xie et al., 2022), DeepREL (Deng et al., 2023), and TitanFuzz (Deng et al., 2024). We have followed the instructions in their source code and repeatedly executed their source code for 48 hours in the same environment to separately detect bugs on TensorFlow and PyTorch frameworks. Note that the final outputs of existing testing methods were thousands of bug candidates, which requires dozens of man-months to validate these candidates, and how to automatically validate and filter out invalid bug candidates is out of our scope. Therefore, we directly recorded ❶ the API coverage and ❷ the number of bug candidates generated by each method in experiments. More details of the experiment results and settings are in our repository (our, 2024). The experimental results are shown in Fig. 9, where the X-axis represents time and the Y-axis represents API coverage and the number of generated bug candidates. In experiments, Muffin covered the fewest APIs (52 TensorFlow APIs) and generated 233 bug candidates within 48 hours. TitanFuzz achieved the largest API coverage with a total of 3,413 APIs on two frameworks. Note that, FreeFuzz and DeepREL have generated a large number of bug candidates, with 737,070 and 379,478 for the two frameworks respectively, which is 49.67 times and 25.57 times the detected buggy candidates of TitanFuzz. However, we have not found an efficient way to automatically verify these bug candidates. So many bug candidates require a significant amount of manual verification and screening, which poses a challenge for comparing the testing effects of different testing methods in real-world scenarios.
@@ 6.1.3. Other Testing on DL Compiler
Comparison Experiment and Results: We have compared three advanced open-sourced DL compiler testing methods to study their test effects, namely Tzer (Liu et al., 2022a), Muffin-TVM and GenCog (Liu et al., 2023c), where Muffin-TVM is a reimplementation of Muffin (Gu et al., 2022a) by researchers, aiming to detect status bugs by converting DL models to relay IR with TVM’s Keras frontend. We have followed their recommendation to repeatedly execute their source code for 4 hours in the same environment to separately detect bugs on TVM (Liu et al., 2022a). Similar to the setting in § 5.2, in the experiments, we collected ❶ the code line coverage and ❷ the number of bug candidates generated by each method. More details are in our repository (our, 2024) The experimental results are shown in Fig. 10, where the X-axis represents time and the Y-axis represents Line coverage and the number of bug candidates. We can observe that Tzer has the lowest line coverage, covering only 26.173 lines after four hours of execution, while the other two methods separately covered 45,610 lines and 53,309 lines within the first hour. Particularly, Muffin-TVM covered more than 50,000 lines of TVM code and generated 10436 bug candidates in the experiment, significantly higher than the results of the other two methods. Our analysis showed that such results could be related to its test case generation methods. Muffin-TVM built diverse DL models using advanced framework testing methods and converted them to TVM for testing. The model graphs carried rich information (e.g., operators) and conformed to the compiler’s semantical constraints, thus effectively exploring DL compiler code lines and discovering bug candidates.

### v3

# Deep Learning Library Testing: Definition, Methods and Challenges

@@ 1. Introduction
However, existing surveys on the DL library testing are limited. We have compared related surveys in Table 1 and observed that existing works mainly focus on testing and repairing DL models or software built upon the models (Braiek and Khomh, 2020; Gezici and Tarhan, 2022; Martínez-Fernández et al., 2022; Zhang et al., 2022; Hu et al., 2023) (italic columns in Table 1). Even though some works involve testing DL libraries, they typically only cover certain components of various libraries and fail to provide a detailed introduction, analysis, and discussion across testing methods for different DL library components. For example, Zhang et al. (Zhang et al., 2020a) and Ji et al. (Ji et al., 2023) discussed testing methods for DL frameworks, while Li et al. (Li et al., 2020b) focused on the workflow and performance of DL compilers. To fill this gap, this paper focuses on the bugs that affect the functionality correctness and efficiency of DL libraries and comprehensively summarizes corresponding testing methods for three DL library components, namely ❶ DL frameworks that execute the DL program and construct DL model, ❷ DL compilers that compile and translate DL models into optimized operators, and ❸ DL hardware libraries that map operators to the hardware to perform calculations. This paper further analyzes the advantages and limitations of existing methods and delves into the challenges and future research opportunities in DL library testing. Our goal is to provide a set of practical findings to promote the development of DL library testing research, thereby ensuring the security and reliability of DL systems built on these libraries in real-world scenarios. Note that the scope of this paper includes the testing methods on three DL library components, but does not include the DL model and DL program testing. The main contribution can be summarized as follows:
  • We present the first comprehensive and detailed survey of testing methods for various libraries in the DL workflow, including DL frameworks, compilers, and underlying hardware libraries. Our work expands and enhances existing DL testing surveys, which only focus on the specific DL component (e.g., DL framework) and lack an in-depth analysis and discussion of bugs and testing methods for different DL library components.
@@ 2.2. DL Library Components
DL libraries implement the abstract DL model based on the program and perform specific operations and optimizations on the underlying hardware to obtain computation results of DL model training and inference. Executing the DL program and building a DL model on the DL underlying libraries mainly involves three components, namely the DL framework, DL compiler, and DL hardware library. The three components are closely related, cooperate with each other, and together form the DL workflow, as shown in Fig. 1. The code in dotted boxes presents the demo inputs of each component. Developers first need to call the DL framework APIs to construct a DL program. As shown in Fig. 1 (a), the DL program calls the DL library APIs to define the convolutional layer (torch.nn.Conv2d), linear layer (torch.nn.Linear), and reshape operation (torch.reshape). DL framework then executes the DL program and constructs the corresponding DL model. Fig. 1 (b) shows the corresponding DL model graph. Subsequently, the DL compiler translates the input model into intermediate representations (IRs) and performs optimization, and outputs the optimized operators and code. During the optimization process, the DL compiler eliminates the ‘Reshape’ in the gray dashed box of Fig. 1 (b) due to functional redundancy. It also reformulates ‘Conv2d’ as the matrix multiplication operator ‘cublasSgemm’ in Fig. 1 (c). Finally, the DL hardware library (e.g., cuDNN) accepts the output of the compiler and maps calculations to the DL hardware to perform calculations and obtain the results.
@@ 3. Research Method
RQ 4.2 What are the challenges in existing DL library testing?
@@ 3.2. Conducting
The search string consists of two major components connected by an ‘OR’ operator. The first component focuses on testing techniques for DL libraries, which is the primary scope of our survey. We use ‘Test OR Testing’ as primary keywords since they effectively capture various testing methodologies, including metamorphic testing, mutation testing, and differential testing. We specifically add ‘Fuzzing’ as an additional keyword because some fuzzing-specific works might not be explicitly labeled with general testing terms. Then we use ‘AND’ to connect it with keywords such as ‘DL Library’ and ‘DL Compiler’ to ensure that these testing methods are relevant to DL library testing. The second component, ‘Deep Learning Library Bugs’, ensures the inclusion of broader empirical studies and bug analysis papers that might not explicitly mention testing methodologies. Through this carefully designed search strategy, we ensure comprehensive coverage of both testing techniques specific to DL libraries and relevant empirical studies, thereby establishing a solid foundation for our systematic literature review.
@@ 5.1.2. Differential Testing on DL Framework
Model-level Differential Testing. The model-level differential testing usually leverages the different results of a widely-used DL model (e.g., ResNet-50) on different platforms or frameworks to detect bugs (Wang et al., 2020b; Li et al., 2022a). CRADLE (Pham et al., 2019) is one of the first tools to detect and identify bugs based on the concept of differential testing. Based on Keras (Ketkar and Ketkar, 2017) which can build and train models on different DL frameworks as backends, CRADLE conducted differential testing on three frameworks(i.e., TensorFlow, CNTK, and Theano) and finally detected 12 bugs. Fig. 7 shows two trigger figures of the inconsistencies that cause one model to have different prediction results and accuracy on different DL frameworks. CRADLE compared the model layer outputs between multiple DL frameworks and detected such inconsistencies. However, the inconsistent outputs of one layer may further lead to inconsistencies in subsequent layers, therefore this localization method is prone to false positives (FPs) and false negatives (FNs). To break these limitations, Guo et al. (Guo et al., 2020) mutated the parameters of model layers to explore more DL model behaviors and leveraged a causal-testing-based technique to localize buggy layers and reduce the FPs. Although the above testing methods effectively detected bugs in popular DL frameworks like TensorFlow, Theano, and CNTK, they only focused on the model inference process and could not detect bugs in the backward propagation stage. To fill the research gap, Gu et al. (Gu et al., 2022a) proposed Muffin, which creatively generated structure information and layer information of the DL model to thoroughly explore possible abnormal behaviors of the model. Then it conducted differential testing between DL frameworks for model training and inference processes and finally detected 39 new bugs. Although model-level differential testing methods obtain outstanding test results, they still have limitations in practice. The major limitation is that they can only explore limited DL model-related APIs and implementations in the framework For example, existing research (Wei et al., 2022) indicated that LEMON (Wang et al., 2020b) only covered 35 TensorFlow APIs. As a result, these methods cannot comprehensively test the entire framework (including APIs that can be used to develop ML algorithms), limiting the effectiveness of model-level differential testing methods.
Fig. 7. Trigger Inputs for Inconsistencies Between DL Frameworks
@@ 5.1.4. Metamorphic Testing on DL Framework
Metamorphic testing has been widely used to detect problems in DL models and systems (Zhang et al., 2021c; Arrieta, 2022). Despite the success of metamorphic testing in DL testing, researchers have paid limited attention to the metamorphic testing of DL frameworks. Ding et al. (Ding et al., 2017) constructed 11 MRS and validated AlexNet on the Caffe DL Framework. Although they were the first to try to design MRs to validate a DL framework, they focused on the DL model accuracy in tests and paid less attention to DL framework bugs. To address the limitation of localizing DL framework bugs at the parameter and tensor level, Chen et al. (Chen et al., 2024) proposed a testing framework with 18 MRs to test 10 common DL operators. However, their manually designed MRs only cover popular DL operators like Conv2d and lack support for the thousands of other DL framework APIs.
@@ 5.1.5. Other Testing on DL Framework
Comparison Experiment and Results. To compare the effectiveness of testing methods based on different testing techniques and principles, we conducted experiment with five state-of-the-art and representative open-source testing methods, covering three popular testing techniques and both model-level and API-level methods, namely Muffin (Gu et al., 2022a), FreeFuzz (Wei et al., 2022), DocTer with conforming (CI) and violating inputs (VI) (Xie et al., 2022), DeepREL (Deng et al., 2023), and TitanFuzz (Deng et al., 2024). We have followed the instructions in their source code and repeatedly executed their source code for 48 hours in the same environment to separately detect bugs on TensorFlow and PyTorch frameworks. Note that the final outputs of existing testing methods were thousands of bug candidates, which requires dozens of man-months to validate these candidates, and how to validate and filter out invalid bug candidates automatically is out of our scope. Therefore, we directly recorded ❶ the API coverage and ❷ the number of bug candidates generated by each method in experiments. More details of the experiment results and settings are in our repository (our, 2024). The experimental results are shown in Fig. 9, where the X-axis represents time and the Y-axis represents API coverage and the number of generated bug candidates in scientific notation. In experiments, Muffin covered the fewest APIs (52 TensorFlow APIs) and generated 233 bug candidates within 48 hours. TitanFuzz achieved the largest API coverage with a total of 3,413 APIs on two frameworks. Note that, FreeFuzz and DeepREL have generated a large number of bug candidates, with 737,070 and 379,478 for the two frameworks respectively, which is 49.67 times and 25.57 times the detected buggy candidates of TitanFuzz. However, we have not found an efficient way to automatically verify these bug candidates. So many bug candidates require a significant amount of manual verification and screening, which poses a challenge for comparing the testing effects of different testing methods in real-world scenarios.
@@ 6.1.3. Other Testing on DL Compiler
Comparison Experiment and Results: We have compared three advanced open-sourced DL compiler testing methods to study their test effects, namely Tzer (Liu et al., 2022a), Muffin-TVM and GenCog (Liu et al., 2023c), where Muffin-TVM is a reimplementation of Muffin (Gu et al., 2022a) by researchers, aiming to detect status bugs by converting DL models to relay IR with TVM’s Keras frontend. We have followed their recommendation to repeatedly execute their source code for 4 hours in the same environment to separately detect bugs on TVM (Liu et al., 2022a). Similar to the setting in § 5.2, in the experiments, we collected ❶ the code line coverage and ❷ the number of bug candidates generated by each method. More details are in our repository (our, 2024) The experimental results are shown in Fig. 10, where the X-axis represents time and the Y-axis represents line coverage and the number of bug candidates in scientific notation. We can observe that Tzer has the lowest line coverage, covering only 26.173 lines after four hours of execution, while the other two methods separately covered 45,610 lines and 53,309 lines within the first hour. Particularly, Muffin-TVM covered more than 50,000 lines of TVM code and generated 10436 bug candidates in the experiment, significantly higher than the results of the other two methods. Our analysis showed that such results could be related to its test case generation methods. Muffin-TVM built diverse DL models using advanced framework testing methods and converted them to TVM for testing. The model graphs carried rich information (e.g., operators) and conformed to the compiler’s semantical constraints, thus effectively exploring DL compiler code lines and discovering bug candidates.
@@ 9. Conclusion
Acknowledgements.
The authors would like to thank the anonymous reviewers for their insightful comments and valuable suggestions. This work is supported partially by the National Key Research and Development Program of China (2023YFB3107400), the National Natural Science Foundation of China (62006181, 62132011, 62161160337, 62206217, U20A20177, U21B2018), and the Shaanxi Province Key Industry Innovation Program (2021ZDLGY01-02 and 2023-ZDLGY-38). Thanks to the New Cornerstone Science Foundation and the Xplorer Prize.